OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle -2x + y = 1 }\) \(\large\color{black}{ \displaystyle -4x + y = -1}\) instead of subtracting them, which might be inconvenient because of sign-confusion, I would multiply one of the equations (which one, is up to you) times (-1). THEN add the equations together.

OpenStudy (anonymous):

I'm not sure about this one because I got, (-1,3)

OpenStudy (solomonzelman):

may I ask you how did you get that? I am not necessarily saying you are wrong, just want to see your work.

OpenStudy (anonymous):

I subtracted 4x from both sides, oh... I did it wrong

OpenStudy (solomonzelman):

alright, you wanted to do a substitution (for y) ? well... elimination would be better here....

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle -2x + y = 1 }\) \(\large\color{black}{ \displaystyle -4x + y = -1 }\) multiply the second equation times -1, and re-write the system after that. \(\large\color{black}{ \displaystyle -2x + y = 1 }\) \(\large\color{black}{ \displaystyle 4x - y = 1 }\) now add the equations together.

OpenStudy (anonymous):

I don't get how that's going to give me the answer, can you tell me what I have to do first?

OpenStudy (solomonzelman):

I am giving you all directions.

OpenStudy (solomonzelman):

do you see "what" i did so far ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

then tell me what is `(y) + (-y) ` equal to ?

OpenStudy (anonymous):

0

OpenStudy (solomonzelman):

y + - y = ?

OpenStudy (solomonzelman):

yes 0

OpenStudy (solomonzelman):

so, when you add the equations (ADD the X's , the Y's the numbers on the right sides of the equations, SEPARATELY). -2x + y = 1 4x - y = 1

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

( -2x + y = 1) \(\normalsize\color{black}{ \rm ^{^{\LARGE +}} }\) ( 4x - y = 1) \(\large \color{black}{ \bf ^{ ^\text{___________________________} } }\)

OpenStudy (anonymous):

6x +0 = 1 ?

OpenStudy (solomonzelman):

close

OpenStudy (solomonzelman):

1 + 1 =? -2x + 4x =?

OpenStudy (solomonzelman):

you got 0 for middle term correctly.

OpenStudy (anonymous):

1 + 1 = 2 and -2x + 4x = 2x since it is a negative 2 in the beginning

OpenStudy (anonymous):

or is it -6x

OpenStudy (solomonzelman):

yes the result of addition of the 2 equations is 2x + 0 = 2

OpenStudy (solomonzelman):

so you can already find the x. x= ?

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

x=1 and knowing that x=1, you can plug in 1 instead of x, into any of the initial (original) equations to solve for y.

OpenStudy (anonymous):

I got y = 3

OpenStudy (anonymous):

So is is (1,3) ? That's what I got. When I plugged in the 1 in the equations, I got 3. Is that right ?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

x=1 y=3 the solution is (1,3)

OpenStudy (solomonzelman):

sorry for late replies, I am lagging

OpenStudy (anonymous):

It's fine and thanks so much!

OpenStudy (solomonzelman):

You are surely welcome!